Skip to Search

 
diagonal spacer

Systems Staff Meeting Notes
January 4, 2008

Present: Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Amy Polk, Mark Sullivan, Stephen Williams
Announcements
  1. Phone problems - the message indicator is not working on the main phone. It is now sending email notifications of voicemail messages to SysHelp, until the problem can be resolved. Waiting to speak with Rosa from Telecom, who will return on Monday.
  2. Dean Russell to revisit the Systems meeting sometime this month.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Web server parts to be ordered.
  2. Systems training - No schedule yet. How about next Thursday? We'll have 2-3 people train with Logan on the SAN and then they will, in turn, train 2-3 people.
  3. Backup Exec - video card died. It is back up now.
  4. Education failover server is dead. Keeps rebooting itself. Carol reported that the UPS is making a noise this morning. Need to replace the power supply.
  5. Web server - major problems. We have it on the virtual server with a version current to December 2007. Bill suggests being cautious bringing it online. The controllers will be ordered for the old server. If that doesn't fix the problem, we can bring it up on the virtual server. We need to register the DLL. We'd prefer to test the virtual server thoroughly before bringing it up.
Networking
  1. McAfee - still nothing.
  2. SAN - it is up here in East and in MSL. DLC mirroring of data is working. All data on the production SAN is backed up. Aliasing is not up yet. There is still some final configuration left to do.
  3. SMS upgrade - Amy has been working on the log-on scripts for the SAN. She will move the SMS up on her list.
  4. Campus AD - Amy had a meeting with Stacey on Monday - figuring out the public AD/print demo. It will be set up in West.
  5. Consolidated logging - nothing new.
  6. Windows upgrade slipstreaming - Cynthia created a CD, but it won't install anything past Service Pack 2. Amy has some further instructions on the procedure.
Software Projects Status
  1. Draft of the blog policy - waiting for Will to review and then send it to the Directors for approval.
  2. Grover and DLC Tracking on hold this week.
  3. Greenstone - new loader/pre-loader working fine.
  4. Finished the Circulation Analysis for Blake and Carrie. Expansion is under discussion.
  5. Winston is tracking down a PeopleSoft problem - terminated employees continue to show up. We are coming up with a list of those who should not show. Winston is working with Tina and Jennifer Ellenwood.
DLC Issues
  1. Nothing new, but Logan should be able to address things soon. He may be able to do the storage upgrade by the end of the month.
Web Projects
  1. Cross over to the new server - editors will need to be reset to point to the new server. The next couple of weeks are a really busy time of year for Web changes due to the start of the new semester.
  2. Stephen to assist Tom with the blog.
Systems Security Management
  1. Meeting held.

 

University of Florida Home Page